OpenStudy (anonymous):

let H(x,y,z) = x^2+y^2+2z^2, and S be the level surface H(x,y,z)= 4. find the coordinates of a point P on the surface S where the tangent plane to S is parallel to the plane 2x+4z=0.

OpenStudy (anonymous):

ok I know that if i find the gradH = normal vector for tangent plane

OpenStudy (anonymous):

\[gradH=2 x i +2yj + 4zk\]

OpenStudy (anonymous):

and I also know that the planes normal is 2i + 4k

OpenStudy (anonymous):

but I don't know whats next can you help me out @eliassaab

OpenStudy (anonymous):

@Jhannybean can you help me please :)

OpenStudy (anonymous):

@Hero can you help me please

OpenStudy (anonymous):

@SolomonZelman can you help me please

OpenStudy (anonymous):

@ganeshie8 can you help me with this one

ganeshie8 (ganeshie8):

heyy

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

this is the question :/

ganeshie8 (ganeshie8):

You're done !

OpenStudy (anonymous):

what do you mean?

ganeshie8 (ganeshie8):

vectors are parallel if the components are proportional

ganeshie8 (ganeshie8):

two planes are parallel if their normal vectors are parallel, yes ?

OpenStudy (anonymous):

so what are the coordinates

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i just know that the y coordinate is 0 but i don't know about x and z

ganeshie8 (ganeshie8):

normal of tangent plane to S : \(2 x i +2yj + 4zk\) you want this parallel to the vector : \(2i + 0j + 4k\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so take the cross product?

ganeshie8 (ganeshie8):

just set the components equal toeach other

OpenStudy (anonymous):

2i +4k = 2xi +2yj + 4zk ?

OpenStudy (anonymous):

or 2i +4k = 4

ganeshie8 (ganeshie8):

sorry was on phone

OpenStudy (anonymous):

its ok dont worry about it

ganeshie8 (ganeshie8):

i mean set the ratio of components equal to each other : \[\dfrac{2x}{2} = \dfrac{4z}{4}\]

ganeshie8 (ganeshie8):

that gives \(x = z\)

ganeshie8 (ganeshie8):

and you know \(y = 0\)

ganeshie8 (ganeshie8):

use the equation of surface to solve for the points

OpenStudy (anonymous):

so x^2 + 2(x)^2 = 4

ganeshie8 (ganeshie8):

yes!

OpenStudy (anonymous):

so x = z = squareroot(4/3)

ganeshie8 (ganeshie8):

Yep!

ganeshie8 (ganeshie8):

wait you should get two points

OpenStudy (anonymous):

negative and positive

ganeshie8 (ganeshie8):

\[x = z = \pm \sqrt{4/3} = \pm 2/\sqrt{3}\] right ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how can i know what point to pick

ganeshie8 (ganeshie8):

so the points would be : \[(2/\sqrt{3}, ~~0,~~ 2/\sqrt{3})\] \[(-2/\sqrt{3}, ~~0,~~ -2/\sqrt{3})\]

OpenStudy (anonymous):

so both

ganeshie8 (ganeshie8):

there can be more than one tangent plane parallel to the given plane

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

yeah thats right

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

thanks I see it